### v1

# UEMM-Air: A Synthetic Multi-modal Dataset for Unmanned Aerial Vehicle Object Detection

Abstract
The development of multi-modal object detection for Unmanned Aerial Vehicles (UAVs) typically relies on a large amount of pixel-aligned multi-modal image data. However, existing datasets face challenges such as limited modalities, high construction costs, and imprecise annotations. To this end, we propose a synthetic multi-modal UAV-based object detection dataset, UEMM-Air. Specially, we simulate various UAV flight scenarios and object types using the Unreal Engine (UE). Then we design the UAV’s flight logic to automatically collect data from different scenarios, perspectives, and altitudes. Finally, we propose a novel heuristic automatic annotation algorithm to generate accurate object detection labels. In total, our UEMM-Air consists of 20k pairs of images with 5 modalities and precise annotations. Moreover, we conduct numerous experiments and establish new benchmark results on our dataset. We found that models pre-trained on UEMM-Air exhibit better performance on downstream tasks compared to other similar datasets. The dataset is publicly available (https://github.com/1e12Leon/UEMM-Air) to support the research of multi-modal UAV object detection models.
1 Introduction
With the advancement of Unmanned Aerial Vehicles (UAV) technology [23, 31]and deep learning [15, 30], vision tasks of UAV have shown great potential in many fields such as agriculture and rescue [31, 33]. The object detection from the perspective of UAV (UAV-OD) [22, 37, 42] has attracted widespread attention from research community. This is due to its ability to efficiently complete tasks such as urban monitoring [39] and military reconnaissance [18]. Unlike general object detection [43], UAV-OD tasks exhibit characteristics such as complex backgrounds and small objects. Therefore, models trained on general object detection datasets [17, 8] can hardly be applied to UAV-OD tasks.
To this end, many scholars have constructed object detection datasets from the perspective of UAV. For example, VisDrone [5] and UAVDT [4] consider various scenes, weather conditions, and environments, providing a good benchmark for UAV-OD tasks. However, with the development of multi-modal learning [1, 38], these datasets are facing challenges such as being single-modality and having insufficient data. DroneVehicle [35] utilizes two image modalities: infrared and visible, with the infrared modality enhancing detection accuracy in nighttime scenes. However, due to the request for manual labeling and aligning two modalities images, the dataset annotation cost is relatively high. AU-AIR dataset [3] takes into account the potential value of UAV parameters for object detection. However, it covers relatively only a few scenes and images have some imprecise annotations. These drawbacks are not conducive to model training.
As illustrated in Fig.1, to address the aforementioned issues, we construct a new synthetic multi-modal UAV object detection dataset UEMM-Air (Unreal Engine Multi-modal Dataset for UAV-based Object Detection). Specifically, we first utilize the Unreal Engine (UE) [7] and AirSim [29] framework to build various simulated scenarios for UAV flights. Subsequently, we implement automatic UAV flight control and collect data at different altitudes, scenes, and modalities. Finally, we design an annotation algorithm to automatically generate object detection labels. It is worth noting that our dataset contains five modalities, including visible, depth, segmentation, surface normals, and UAV IMU parameters. We also provide more accurate annotation information, including fine-grained and scene labels, which are beneficial for various types of object detection tasks. Additionally, we conduct experiments on multiple types of tasks in the field of object detection, demonstrating the research significance of our dataset across various tasks. The main contributions are as follows:
We propose a synthetic multi-modal UAV-based object detection dataset, UEMM-Air. It includes a greater number of modalities, more accurate annotations, richer scenes, and fine-grained labels than existing datasets (e.g. SynDrone). To the best of our knowledge, UEMM-Air is the UAV-OD dataset with the largest number of modalities.
We introduce a new heuristic algorithm for automatic data annotation. Compared with labeling strategies in SynDrone, ours can provide more accurate annotations by introducing segmentation and depth modalities to enhance the identification of objects. Especially when it comes to addressing visually overlapped objects.
We conduct experiments on multiple types of tasks in the field of object detection, providing new benchmark results. Due to our accurate labeling and rich modalities, our UEMM-Air provides remarkable transferability for downstream tasks.
| Dataset | Year | MM | # modalities | # images | # classes | Size [px] | Angle |
| Stanford-Drone [28] | 2016 | ✗ | - | - | 7 | 1450×1080 | 90 |
| UAVDT [4] | 2018 | ✗ | - | 40376 | 3 | 1080×540 | variable |
| VisDrone [5] | 2018 | ✗ | - | 8629 | 10 | 1920×1080 | variable |
| AU-AIR [3] | 2020 | ✓ | 2 | 32823 | 8 | 1920×1080 | 45 to 90 |
| Drone-Vehicle [34] | 2022 | ✓ | 2 | 28 | 5 | 640×512 | 90 |
| SynDrone [27] | 2023 | ✓ | 3 | 60 | 9 | 1920×1080 | 30,60,90 |
| UEMM-Air (Ours) | 2024 | ✓ | 5 | 20 | 13 | 1920×1080 | variable |
2 Related Work
2.1 UAV-based Object Detection
Due to the typically top-down perspective of UAVs and the relatively small scale of the objects, general object detection methods [9, 25, 24, 2, 20] tend to be not suitable for UAV-based Object Detection (UAV-OD) tasks. The mainstream methods for UAV-based object detection primarily employ coarse-to-fine strategies [6, 16, 40]. Initially, the detection process focuses on identifying larger objects, while concurrently pinpointing dense subregions containing small objects. These subregions are subsequently utilized as inputs for the model to refine detection results. For example, a CZDetector [21]employed a density-based cropping algorithm to identify regions with crowded objects and then increased the size of those regions to enhance the training dataset. Alternatively, Koyun et al. [14] utilizes a Gaussian mixture model to supervise the detector in generating object clusters composed of focusing regions. And DMNet [16] optimizes region selection by conducting a density map guided connected crop generation.
2.2 UAV-OD Datasets
The UAV-based object detection dataset provides multi-class images and videos captured by UAVs. These datasets are significant for promoting the research and development of various computer vision tasks, including object tracking, path planning, and scene understanding. We summarized several commonly used UAV-OD datasets in Table 1. Due to the high cost of obtaining images taken by UAV, early datasets, such as UAVDT and VisDrone, have only a single modality, namely RGB images. In addition, owing to hardware limitations, datasets like Drone-Vehicle are small, low-resolution, and difficult to expand.
Stanford-Drone [28] is a large-scale dataset containing overhead images and videos of multi-class of objects moving and interacting at Stanford University. This dataset can be used for learning and evaluating multi-object tracking, activity understanding, and trajectory prediction.
UAVDT [4] has 80,000 representative frames which are annotated with bounding boxes and 14 kinds of attributes in various complex scenarios. It focuses on 3 specific computer vision tasks: object detection, single-object tracking, and multiple-object detection.
VisDrone [5] is a large-scale benchmark dataset in object detection and tracking with various environment conditions and camera viewpoints. It contains 10 categories objects of frequent interest in drone applications and more than 2.5 million annotation bounding boxes.
AU-AIR [3] includes extracted frames meta-data, bounding box annotations for traffic-related object categories, and multi-modal flight sensor data. The dataset is captured at low altitudes at the intersection and covers various lighting conditions.
Drone-Vehicle [34] offers a drone-based RGB-Infrared cross-modality vehicle detection dataset and corresponding precise annotations. This dataset covers multiple scenarios and objects from day to night with three different angles and heights.
SynDrone [27] proposes a multi-modal synthetic benchmark dataset containing both images and 3D data taken at multiple flying heights. It includes 28 classes of pixel-level labeling and object-level annotations for semantic segmentation and object detection.
3 UEMM-Air Dataset
3.1 Scene Construction and Flight Control Logic
Previous UAV-OD datasets have been limited in scene diversity, which tends to affect model generalization. Therefore, we aim to construct a dataset with richer scenes to improve the performance of the models. To be specific, we utilize Unreal Engine with CityBLD [32] plugin. It can create cities of almost any size and style in a very short time to simulate scenarios in the real world. we build several scenes in Unreal Engine, including cities, parks, highways, etc. We collect a total of 13 categories and more than hundreds of vehicle models.
We leverage Unreal Engine’s movement animation to simulate dynamic scenes in reality. Employing the traffic features of Unreal Engine, we can flexibly design and construct various complex road layouts, including city streets, highways, and country roads. These layouts can precisely simulate real-world terrain and traffic conditions, providing realistic infrastructure for game scenes. Additionally, we can generate a wide variety of vehicles in the virtual environment. These vehicles can automatically navigate the generated roads based on predefined traffic rules and behaviors. By setting paths and control parameters, the vehicles can simulate real traffic flow, obey traffic signals, avoid pedestrians, and respond to traffic congestion, thereby creating highly realistic dynamic traffic scenarios.
To collect data, we controll the UAV to fly and take pictures in Unreal Engine. Specifically, we build an Unmanned Aerial Vehicle simulator using AirSim and Pygame. When flying to a satisfactory shooting point, we control the UAV to fly within a height range of 5 meters to 50 meters above the horizontal surface, taking a set of photos every 5 meters up. The camera rotates from 0 degrees to 90 degrees by 5 degrees for each step. To obtain aligned pictures of different modalities, our simulator temporarily stops running when taking photos.
3.2 Automatic Image Segmentation and Annotation
Most of the existing object detection datasets are manually annotated. Manual image annotation faces challenges in terms of accuracy and efficiency, especially when dealing with a large number of labels or low-resolution images.
To avoid manual annotation, the SynDrone dataset employs an automatic image labeling algorithm. They derive the absolute coordinates of UAV and vehicles from Unreal Engine, then obtain the detection frame of the object by analyzing their relative position. However, this strategy causes some incorrect annotations where objects are visibly blocked but their coordinates are still marked on the image, as illustrated in Fig.2.
In order to alleviate the problem of mislabeling in the SynDrone dataset, we propose a heuristic automatic image annotation algorithm. It makes full use of semantic and distance information from segmentation and depth images to avoid labeling visually blocked objects and mislabeling overlapped ones. Our approach is illustrated in Fig. 3.
Employing the AirSim simulator, we assign the same color label to the same class of objects in the Unreal Engine environment. For each class, we convert contour detection on objects into bounding boxes and get the initial annotation. However, this step cannot recognize objects of the same category that are overlapped in the segmentation image and will mark them as one object.
To avoid mislabeling visually overlapped objects, we utilize depth images where pixel value represents the distance from the object to the camera plane to perform a secondary annotation. Intuitively, depth values mildly change on each object and a depth value jump indicates multiple objects existences. Therefore overlapped objects can be correctly identified through depth observation. We detect depth mutations within segmented bounding boxes to confirm object edges, adjusting labels accordingly.
Fig. 4 presents sample annotation results for comparing our proposed algorithm with only utilizing segmentation information. It can be observed from Fig. 4 (a) that leveraging segmentation information alone can’t effectively handle cases of visual overlapped (as shown by the pink box). Our approach can alleviate this issue by correctly distinguishing the two vehicles within the pink box. Fig. 4 (b) shows the numerical statistics of the annotations generated by the two methods. It can be observed that our method successfully annotates more objects, because our approach can distinguish overlapped instances and correct the annotations accurately employing depth information.
3.3 Acquisition Setup
We adopt our camera sensor setup for the AirSim simulator to ensure diversity in data. The acquisition pipeline equips the UAV with several co-registered sensors. With the help of these sensors, we collect 5 modalities: RGB, infrared, segmentation, surface normal, and IMU parameters.
RGB Camera: It offers a resolution of pixels. The vertical field of view (FoV) increases dynamically from to , indicating that the viewing angle changes from a horizontal to a top-down view. All RGB images are stored in PNG format.
RGB images contain rich color, and spatial information, facilitating better image understanding and object recognition. The visual image is the most common modality in computer vision tasks. However, in complex environments such as nighttime, visual images alone may not perform well due to the poor visibility and the resulting inability to effectively detect objects.
Depth Camera: The depth camera has the same FoV, resolution and storage format as the RGB camera. It interpolates each pixel value from 0 to 255 according to the depth of the distance from the camera plane. The white pixels show a depth of more than 100 meters, while the black pixels indicate a depth of 0 meters.
Depth images leverage pixels to represent the distance from the object to the camera, reflecting the spatial shape and structure of the photographed scene. Therefore, we utilize it to address the issue of inaccurate annotations caused by overlapped visual information of the objects during the annotation generation process. It can also be leveraged to deduce an object’s height, convexity, and relative position, which aids in multi-modal object objection.
Segmentation Camera: The segmentation camera maintains the same FoV, resolution, and data format as the preceding cameras. It generates distinct colors for pixels belonging to different categories of objects to ensure accurate segmentation of the scene.
The segmentation image divides the image into multiple regions with similar attributes, providing pixel-level information where each pixel is assigned to a precise category label. Because of the detailed segmentation information, this modality can assist in the automatic generation of detection annotations. Additionally, since segmentation images inherently contain positional information, combining them with other modalities for detection often leads to improved accuracy.
Surface Normal Camera: The Surface Normal Camera maps the X, Y, and Z components of the surface normal to an RGB range from 0 to 255. Due to the gradual changes in normal direction, it is difficult to distinguish. Therefore, the contrast of the normal camera images is set as 1.5 to more distinctly delineate changes in the direction of the normals. This camera saves its pictures in PNG format and has the same field of view and resolution as the Scene Camera.
Surface normal images primarily capture the geometric features and surface details of the target object. When fused with RGB images or other modalities, they can compensate for deficiencies in texture features. For example, in fine-grained object detection, the texture features introduced by the surface normal modality can help the model learn deeper fine-grained information. Additionally, it reveals intricate surface details essential for generating precise 3D reconstruction surface models.
IMU Parameters: IMU parameters encompass dynamic state data, GPS information, flying altitude and timestamps. The dynamic state parameters consist of attitude angle, linear velocity, body angular velocity, linear acceleration, and collective angular acceleration.
IMU parameters comprise real-time attitude details and UAV position coordinates. In multimodal tasks, the current flying altitude of the UAV can be utilized to assist in determining the scale information of the object. For example, the current frame’s flight posture is beneficial for the model to predict the next frame’s object location, especially in tasks like video object detection or object tracking. Additionally, the UAV’s GPS information can also be employed for post-detection localization tasks.
| Method | Fine-tuned | Pre-trained | ||
| YOLOv8 | VisDrone | SynDrone | 25.6% | 16.0% |
| UEMM-Air | 28.3% | 17.7% | ||
| UAV-DT | SynDrone | 85.4% | 55.6% | |
| UEMM-Air | 86.2% | 56.1% | ||
| FaterRCNN | VisDrone | SynDrone | 5.1% | 1.1% |
| UEMM-Air | 5.5% | 2.3% | ||
| UAV-DT | SynDrone | 48.0% | 9.5% | |
| UEMM-Air | 53.8% | 14.7% |
4 Benchmark and Experiments
4.1 Experimental Setup
Benchmark Models. We adopted YOLOv8-L [12], FasterRCNN [26] with ResNet50 [10], RT-DETR-L [41] as the general object detection baseline models. In the multi-modal object detection experiment, we designed a dual-path multi-modal detector with mid-level feature fusion. We utilized YOLOv7-L [36] as the base detector, with two separate backbone networks to extract features from two modalities. We also designed a feature fusion module that utilizes Coordinate Attention (CA) [11]. Specifically, we first directly concatenated the features of two modalities and employed Coordinate Attention to fuse them simultaneously in terms of channel and spatial information. The fused features were then entered into the neck part of the detector to complete the remaining detection tasks.
| Method | LargeVehicle | MediumVehicle | SmallVehicle | |
| YOLOv8 | 55.6% | 73.7% | 75.9% | 68.4% |
| FasterRCNN | 32.4% | 41.5% | 31.2% | 34.9% |
| RT-DETR | 49.8% | 55.8% | 77.5% | 61.1% |
Training Settings. All experiments were conducted in Pytorch with a NVIDIA RTX 3090 GPU. During the model transferability verification, we set the batch size to 16 and trained for 200 epochs. In other experiments, we froze the backbone network of the detector and trained for 50 epochs with a batch size of 32. All detectors were trained using an Adam optimizer [13] with a momentum of 0.937. The learning rate was initialized as 0.001 with a cosine decay [19]. We fix random seed to 18 to ensure the experiment’s reproducibility.
4.2 Transferability Verification
To demonstrate the advantage of our UEMM-Air on model transferability, We pre-trained two detectors utilizing SynDrone and UEMM-Air, respectively. Then we subsequently fine-tuned the models on the VisDrone and UAVDT datasets. The experimental results are presented in Table 2. While the number of images is smaller than SynDrone (20 & 60), the model pre-trained on the UEMM-Air dataset demonstrates stronger generalization performance on real-world scenario data. For example, after obtaining pre-trained weights on UEMM-Air and SynDrone datasets, we fine-tuned the YOLOv8 model on the VisDrone dataset. The model pre-trained on UEMM-Air demonstrated a 2.7% improvement in and a 1.7% improvement in . This might be attributed to the provision of more accurate annotations, more categories, and more diverse scenarios in UEMM-Air for the model pre-training process.
4.3 Evaluation on Object Detection
In this section, we selected several mainstream detectors and conducted experiments on object detection tasks at coarse-grained and fine-grained labels and multi-modal object detection tasks. Then we conducted analysis based on the model performance and experimental results. These experimental results will serve as the baseline results of our dataset for future research.
| Method | BoxTruck | Bulldozer | Excavator | ForkLift | Jeep | Motor | Pickup |
| YOLOv8 | 61.3% | 62.5% | 42.2% | 51.0% | 72.6% | 75.9% | 78.7% |
| FasterRCNN | 33.9% | 51.5% | 19.4% | 24.4% | 40.6% | 30.7% | 50.6% |
| RT-DETR | 69.8% | 58.2% | 55.9% | 61.2% | 23.1% | 79.8% | 81.8% |
| Method | SUV | Trailer | Truck | Van | RoadRoller | Sedan | |
| YOLOv8 | 81.9% | 51.1% | 56.1% | 69.0% | 56.7% | 71.6% | 63.9% |
| FasterRCNN | 46.1% | 30.6% | 29.5% | 35.9% | 39.6% | 40.1% | 36.4% |
| RT-DETR | 85.1% | 44.3% | 58.5% | 79.2% | 63.2% | 79.4% | 64.6% |
General Object Detection. We employed coarse-grained and fine-grained labels for training three detectors, namely YOLOv8, FasterRCNN, and RT-DETR, respectively. Experimental outcomes are presented in Table 3 and Table 4. In the coarse-grained detection experiments, YOLOv8 achieved the highest precision, with an of 68.4%, followed by RT-DETR with a precision of 61.1%. In the fine-grained detection experiments, RT-DETR exhibited the highest at 64.6%. Across both experiments, FasterRCNN consistently demonstrated lower performance compared to the other two detectors, consistent with the performance of the three detectors in the general domain.
| Modality | ||
| RGB | 68.4% | 53.2% |
| RGB + Seg | 75.3% | 54.1% |
| RGB + Surface | 74.3% | 54.4% |
| RGB + Depth | 73.0% | 54.2% |
Multi-modal Object Detection. In Table 5, we conducted mid-level fusion experiments for multi-modal object detection with RGB modality and other three modalities. The model fusion of RGB with segmentation modality achieved the best performance on , surpassing the baseline model (RGB only) by 6.9%. The fusion of RGB with surface normal modality achieved the best performance on , surpassing the baseline model by 1.2%. However, fusion with depth modalities resulted in the lowest performance. This could be due to the distinct features of object positions in segmentation modality and the detailed texture features in surface normal, both containing more effective information compared to depth.
| Fold | Train | Valid | ||
| 1 | 1,2,3,4 | 5 | 63.9% | 52.6% |
| 2 | 1,3,4,5 | 4 | 62.6% | 53.4% |
| 3 | 1,2,4,5 | 3 | 63.1% | 52.8% |
| 4 | 1,2,3,5 | 2 | 62.3% | 53.7% |
| 5 | 2,3,4,5 | 1 | 64.3% | 53.3% |
4.4 Evaluation on Automatic Annotation Algorithm
Considering that our labeling algorithm is auto-generated, it is necessary to validate the reliability of the labels we generate. In this section, we demonstrated the effectiveness of our labels through cross-validation experiments and visual comparisons with manual labels.
Verification of the Annotations’ Reliability. We randomly divided the dataset into 5 parts and conducted five-fold cross-validation experiments on the YOLOv8 model. The experimental results are demonstrated in Table 6. By sequentially using different sets of 4 parts as the training set and the remaining part as the validation set, we observed that the results of the 5-fold cross-validation were quite similar. The lowest was 62.3%, and the highest was 64.3%, with a range of 2.0%. This result indicates that the annotations we generated are consistent in their distribution.
Comparison with Manual Annotations. We randomly selected a subset of images for manual annotation and then visually compared them with our automatic annotations. As presented in Fig 5, we visualized our generated labels and manually annotated labels separately for comparison. It can be observed that our annotations are almost identical in position to the manual labels. Moreover, our annotation algorithm has some advantages in labeling small objects. We found that when objects are far away, manual annotations may contain errors due to the smaller scale of the objects. For example, manual annotations may not be as closely aligned with the edges of the objects as our generated annotations. It will introduce more foreground information, which could impact the model’s accuracy.
5 Conclusion
In this paper, we released a synthetic UAV-based object detection dataset, named UEMM-Air. Our work achieved three main breakthroughs. Firstly, to the best of our knowledge, UEMM-Air is the UAV-OD dataset with the largest number of aligned modalities (RGB, segmentation, depth, surface normals, and IMU parameters). Secondly, we designed a new automatic annotation method to for UAV-OD tasks, enhancing the accuracy of annotations by employing segmentation and depth images. Finally, we conducted a series of experiments. The results validated that the models pre-trained on UEMM-Air exhibit strong transferability. We also established new benchmarks across various types of UAV-OD tasks, including fine-grained object detection and multi-modal object detection. We will continue to build new simulated scenarios in the future to expand the scale and number of modalities in our dataset, supporting research on UAV vision tasks.
Broader Impacts and Limitations
Broader impacts. Our UEMM-Air dataset features 5 aligned modalities, making it the dataset with the highest number of aligned modalities known to date. This will effectively support UAV multi-modal object detection tasks. Furthermore, leveraging the advantage of easily accessible generative data, the paired multi-modal data we provide will also be beneficial for supporting large-scale multi-modal pre-training tasks.
Limitations. Due to the simulated scenes still requiring manual construction, there are still certain manual costs associated with building the dataset. However, the development of scene dialogue generation tasks in the Unreal Engine is already under research. We believe that it won’t be long before automated scene construction is achieved, integrating with our construction methods to efficiently generate vast amounts of data for model pre-training.
Acknowledge
This work was partially supported by the Fundamental Research Funds for the Central Universities (No. B240201077), National Nature Science Foundation of China (No. 62372155 and No. 62302149), Aeronautical Science Fund (No. 2022Z071108001), Joint Fund of Ministry of Education for Equipment Pre-research (No. 8091B022123), Water Science and Technology Project of Jiangsu Province under grant No. 2021063, Qinglan Project of Jiangsu Province, Postgraduate Research & Practice Innovation Program of Jiangsu Province (No. SJCX24_0183).